A certain flight arrives on time 86 percent of the time. Suppose 160 flights are randomly selected. Use the normal approximation to the binomial to approximate the probability that

​(a) exactly 150 flights are on time.

​(b) at least 150 flights are on time.

​(c) fewer than 133 flights are on time.

​(d) between 133 and 136​, inclusive are on time.
Answer #1

a)

The following information about the Binomial distribution is provided: = - 160 The population proportion of success is p=0.86

b)

0 = Vn.p. Vn.p. (1 – p) = V160 - 0.86 · (1 – 0.86) = 4.3891 Therefore, we get that Pr(X Binomial > 150) Pr(X Normal > 149.5)

c)

Pr(XBinomial < 133) — Pr(X Normal < 133.5) X Normal = Pr 137.6 4.3891 133 – 137.6 4.3891 = Pr | Z < (2 133 – 137.6 4.3891 = P

d)

Pr(133 < X Binomial < 136) Pr(132.5 < X Normal < 136.5) = Pr 132.5 137.6 4.3891 X Normal – 137.6 4.3891 136.5 – 137.6 4.3891

please like ??